OpenStudy (anonymous):

Please Help! Fan & Medal Here is the parent function: y=x^2 Describe how each function below is different from the parent function above ( shifted up 5, right 7, etc)

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ y=x^2 }\) is a parabola at the origin. The center of this one is (7,5) that's all.

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ y=(x-7)^2+5 }\)

OpenStudy (zale101):

y=a(x-h)^2+k h is the horizontal shift to the right by h units. k is the vertical shift up by k units. y=a(x-(-h))^2-k h is the horizontal shift to the left by h units. k is the vertical shift down by k units.

OpenStudy (zale101):

a determines the vertical stretching of the prabola